for each of the following ions, indicate the noble gas that has the same lewis structure as the ion.
express your answer as a chemical symbol?

BR-
O2-
Rb+
Ba2+
Answer #1

1. Br-: Number of electrons in Br- = 35 + 1 = 36. Noble gas that has 36 electrons is Krypton, Kr. So, the noble gas that has the same Lewis structure to Br- is Kr.

2. O2-: Number of electrons in O2- is 8 + 2 = 10. Noble gas that has 10 electrons is Neon, Ne. So, the noble gas that has the same Lewis structure as O2- is Ne.

3. Rb+: Number of electrons in Rb+ = 37 - 1= 36. Noble gas that has 36 electrons is Kr. So, the noble gas that has the same Lewis structure as Rb+ is Kr.

4. Ba2+: Number of electrons in Ba2+ = 56 - 2 = 54. Noble gas that has 54 electrons is Xenon, Xe. So, the noble gas that has the same Lewis structure as Ba2+ is Xe.